OpenStudy (anonymous):

Examine the continuity at x=0

OpenStudy (anonymous):

\[f(x)=\left\{\begin{matrix} \frac{\sqrt{1-cos2x}}{x} & x \ne 0 \\ 0 & x=0 \end{matrix}\right.\]